Can you please help me with this one? ** A restaurant has fixed costs of $137.50 per day and an average unit cost of $4.75 for each meal served. If a typical meal costs $6.00, how many customers must eat at the restaurant each day for the owner to make a profit?
Found 2 solutions by Nate, anjnewman:
Answer by Nate(3500)

there is a profit of $1.25 per meal to the customers
let us take multiples of 1.25
1.25*20=25
1.25*80=100
1.25*100125
1.25*106=132.5
1.25*110=137.5
the store makes a profit when they sell 111 meals to the people!

Answer by anjnewman(1)

137.50 + 4.75x = 6x
subtract 4.75 from 6x = 1.25x
137.50 = 1.25x
divide
137.50 /110x
x = 110